OpenStudy (doodleduck8):

Let g(x) = x^2 + 2x - 6. Find the value of 2[g(2)]. 12 4 18 6

jhonyy9 (jhonyy9):

do you can calcule g(2) = ?

jhonyy9 (jhonyy9):

just substitute the 2 in place of x inside this quadratic

OpenStudy (jdoe0001):

\(\bf g(x) = x^2 + 2x - 6 \\ \quad \\ g({\color{brown}{ 2}})=({\color{brown}{ 2}})^2 +2({\color{brown}{ 2}})-6\implies \square ? \\ \quad \\ 2[g(2)]\implies 2\cdot \square ?\)

OpenStudy (doodleduck8):

I still don't know what the answer is

jhonyy9 (jhonyy9):

exactly how @jdoe0001 ve wrote it above

OpenStudy (doodleduck8):

is it 12 or 4

jhonyy9 (jhonyy9):

first calcule the g(2) = ?

OpenStudy (doodleduck8):

I am not sure how

jhonyy9 (jhonyy9):

substitute in place of x the 2 and make the calculi

OpenStudy (doodleduck8):

so is it 4

jhonyy9 (jhonyy9):

please show your work

OpenStudy (doodleduck8):

g(2), insert 2 for x g(2) = 2^2 + 2(2) + 6 = 4 + 4 + 6 = 4 @jhonyy9

jhonyy9 (jhonyy9):

not is plus 6 is there -6

jhonyy9 (jhonyy9):

4+4-6=8-6=2

OpenStudy (doodleduck8):

so it is 6

jhonyy9 (jhonyy9):

why 6 ?

jhonyy9 (jhonyy9):

g(2)=8-6=2

jhonyy9 (jhonyy9):

so and 2*g(2) = 2*2 = ?

OpenStudy (doodleduck8):

18 is the answer

jhonyy9 (jhonyy9):

why 18 ? 2*2 = ?

jhonyy9 (jhonyy9):

2*2 = not 18

OpenStudy (doodleduck8):

4 is the answer, this is what I said first

jhonyy9 (jhonyy9):

yes but i wan checking your work - ok. ? how you get this 4

jhonyy9 (jhonyy9):

yes 4 is the right answer sure bc. 2*g(2) = 2*2 =4

jhonyy9 (jhonyy9):

hope helped good luck bye bye

OpenStudy (doodleduck8):

Ok, thanks bye